How do I find two numbers that differ by 8 and whose product is 273 by setting up and solving with the quadratic equation?
Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

x - y = 8 ___ x = 8 + y

x * y = 273

substituting ___ (8 + y) * y = 273 ___ y^2 + 8y - 273 = 0

use quadratic formula to find y , then substitute to find x


ANSWER BELOW












factoring ___ (y + 21)(y - 13) = 0

Answer by stanbon(75887)

How do I find two numbers that differ by 8 and whose product is 273 by setting up and solving with the quadratic equation?
---------------------
Equations:
x - y = 8
xy = 273
----
Substitute for x and solve for "y":
(y+8)y = 273
y^2+8y-273 = 0
(y-13)(y+21) = 0
-----
1st Solution:
y = 13 so x = 21
---
2nd Solution:
y = -21 so x = -13
======================
Cheers,
Stan H.